# Nitrogen-Vacancy Centers in Epitaxial Laterally Overgrown Diamond: Towards Up-scaling of Color Center-based Quantum Technologies

### v1

@@ IV Conclusion
The findings in this study indicate low effective stress throughout the top layer of our overgrown material, indicating restrained dislocation propagation from the substrate. The regions overfilling the holes show, on average, a 20-70% reduction of effective crystal strain in comparison to regions directly grown on the heteroepitaxial grown diamond as substrate. We observe an effective decoherence time (reaching 0.5 ms ) comparable to high quality homoexpitaxial CVD grown crystal when decoupling the spin defects from slowly varying noise in the crystal. This characteristic is advantageous for the development of robust diamond-based quantum sensors. The presented study is limited in resolving the complete stress-tensor, and thereby the directional effect of the stress due to dislocations. This can be overcome by studying similar sample containing dense ensemble of NVs that are randomly oriented. Despite low strain, further modification of the growth process might be investigated to also enhance times. The effect of surface polishing, diamond nanofabrication processes, and the possibility of growing preferentially oriented NV ensemble in the single crystal remains to be tested. Our results pave the way towards using large area heteroepitaxial diamonds as a growth substrate for large sized single crystal diamonds for quantum technologies.

### v2

@@ IV Conclusion
The findings in this study indicate low effective stress throughout the top layer of our overgrown material, indicating restrained dislocation propagation from the substrate. The regions overfilling the holes (H regions) show axial (non-axial) stress components ranging between 0.8-2 (2.67-2.94) MHz, in comparison to 2.13-2.90 (3.36-3.58) MHz for regions directly grown on the heteroepitaxial grown diamond as substrate (NH regions), indicating, on average, a 20-70% reduction in effective crystal stress component. We observe an effective decoherence time (reaching 0.5 ms ) comparable to high quality homoexpitaxial CVD grown crystal when decoupling the spin defects from slowly varying noise in the crystal. This characteristic is advantageous for the development of robust diamond-based quantum sensors. The presented study is limited in resolving the complete stress-tensor, and thereby the directional effect of the stress due to dislocations. This can be overcome by studying similar sample containing dense ensemble of NVs that are randomly oriented. Despite low stress, further modification of the growth process might be investigated to also enhance times. Our results pave the way towards using large area heteroepitaxial diamonds as growth substrate for up-scaling single crystal diamonds for quantum technologies: For future sensor device fabrication, centimeter-sized heteroepitaxial substrates with an array of millimeter sized, overgrown holes allow for an effective increase in usable diamond area. Despite the fact that only the diamond over the holes will be optimized for sensor fabrication, usable areas will still be significantly larger than standard homoepitaxial substrates limited to few millimeters size. Further processing steps will be necessary to use these diamond for sensor fabrication, e.g. thinning the samples from the backside to retrieve only the topmost layer (closest to the growth surface). While this is a standard technique for small diamond samples, such processes also need to be tested for larger diamonds. The effect of polishing and thinning to retrieve only the high quality layer, diamond nanofabrication processes, and the possibility of growing preferentially oriented NV ensemble in the single crystal remains to be tested.